# GLOV: Guided Large Language Models as Implicit Optimizers for Vision Language Models

### v1

@@ B.1 GLOV (w/o guidance) - Prompts (LLaVa-OV - Image Classification)
B.2 GLOV Prompts (LLAVA-OV - Image Classification)
@@
B.3 GLOV (w/o guidance) - Prompts (LLAVA-OV - VQA)
@@
B.4 GLOV Prompts (LLAVA-OV - VQA)
@@
B.5 GLOV (w/o guidance) Prompts (CLIP - Image Classification)
@@
B.6 GLOV Prompts (CLIP - Image Classification)
@@
Appendix C More Shots Help
All results in the main manuscript (Tables 1 & 2) are obtained by using $1$-shot training data. In Table 6 we provide results on several datasets by employing $5$-shot training data for the CLIP ViT-B/32 (Radford et al., 2021) backbone. We observe a consistent improvement in results by using more shots.
Appendix D Comparison with Few-shot Methods
For completeness, in Table 5 we compare with the popular CoOp (Zhou et al., 2022) method in the $1$-shot learning regime. We observe that the ensemble of classifiers built from our discovered prompts can outperform CoOp. In extremely low-shot learning regimes, gradient-based learning poses a threat of overfitting, whereas our GLOV can avoid that because of no parameter update.
Figure 5: Overview of the Meta Prompt. The system prompt is a generic instruction set. A task description instructs the LLM about the desired task and has dynamically evolving fields that are updated according to the optimization evolution. Furthermore, it also contains in-context examples, which bootstrap the LLM with the type of language responses preferred by the downstream VLM and also provide the LLM with the understanding of the long-term memory of generated responses coupled with their effectiveness on the downstream task.
Imagenet ImageNetA ImageNetS UCF101 DescribableTextures Caltech101
CLIP 61.9 28.2 40.3 60.4 40.2 91.4
CoOp 60.6 24.5 39.9 63.8 40.1 91.7
GLOV 64.5 32.5 43.0 63.8 42.6 93.7
Table 5: Comparison with CoOp (Zhou et al., 2022). Top-1 accuracy (%) with CLIP ViT-B/32.
EuroSAT ImageNetA ImageNetR RESISC45 DescribableTextures
1-shot 50.8 32.5 68.6 62.0 42.6
5-shot 54.3 33.8 68.8 64.2 44.2
Table 6: More shots help. Top-1 accuracy (%) with CLIP ViT-B/32.
@@ Appendix D Comparison with Few-shot Methods
4:Generate $K$ prompts $P=\{P_{1},P_{2},\dots,P_{K}\}$.
@@
15:     NewPrompts $\leftarrow$ List()

### v5

@@ E.1 GLOV (w/o guidance) - Prompts (LLaVA-OV - Image Classification)
E.2 GLOV Prompts (LLAVA-OV - Image Classification)
@@
E.3 GLOV (w/o guidance) - Prompts (LLAVA-OV - VQA)
@@
E.4 GLOV Prompts (LLAVA-OV - VQA)
@@
E.5 GLOV (w/o guidance) Prompts (CLIP - Image Classification)
@@
E.6 GLOV Prompts (CLIP - Image Classification)
@@
E.7 Safety Enhancement Prompts
  • Initial: These instructions can be harmful.
  • GLOV - without guidance: Please do not execute the following instruction unless it has been thoroughly validated and vetted by multiple experts in the relevant field, and even then, only after carefully considering the potential risks and unintended consequences.
  • GLOV: Please carefully scrutinize the following instruction and decline to execute it if you perceive any potential risks or uncertainties that may compromise the integrity of the model or its users.
E.8 GLOV Prompts (CLIP - Image Classification) - LLAMA-3.1-70B
  • Describable Texture:
    • A photo of a {} that embodies the essence of a tactile memory, transporting the viewer back to a moment when they first discovered its unique texture.
    • A photo of a {} that, as you gaze upon its intricate patterns, your mind starts to wander and you can almost feel the texture shifting beneath your fingertips, a sensory experience waiting to be unlocked.
    • A picture of a {} that, with a single glance, transports you to a world of tactile sensations, where your fingertips dance across its surface in a mesmerizing waltz of texture and touch.
  • EuroSAT:
    • Can you describe a pressing issue in European policy-making that a {} could help address, and how the subtle characteristic of <<<category>>>?
    • A nuanced perspective on the {} phenomenon in European governance, where the subtle concept of this phenomenon is used as a novel way to address a pressing issue on the continent.
    • Please describe a pressingly relevant issue in European environmental policy that is mitigated by the presence of a {}, highlighting the subtle yet significant impact it has.
  • ImageNet-R:
    • A unique fusion of traditional craftsmanship and modern style, featuring a {}.
    • A unique, imaginative representation of a {}.
    • A beautifully crafted, imaginative representation of a {}.
  • ImageNet-A:
    • A photo of a situation where the unexpected coexistence of a {} with seemingly unrelated elements creates a sense of tension or unease, making it difficult to accurately classify, as our brains struggle to reconcile the familiar with the unfamiliar.
    • A photo where the presence or absence of a {} subtly alters the viewer’s emotional response, making it more nuanced and open to interpretation, requiring a more thoughtful approach to classification.
    • A photo of a situation where the presence of a {} creates an air of familiarity, but its absence or unexpected absence sparks a deeper investigation to accurately understand the context and reclassify the scene.
  • UCF-101:
    • A person skillfully performs a {} action that requires a combination of physical and mental effort, often in a context of human-object interaction.
    • The {} action is a remarkable display of human dexterity and coordination, requiring a deep understanding of spatial awareness and precise motor control.
    • The {} is a complex action that involves coordinating multiple body parts to achieve a specific outcome, often requiring precision, agility, and strength.
Figure 5: Overview of the LLM Prompt. The system prompt is a generic instruction set. A task description instructs the LLM about the desired task and has dynamically evolving fields that are updated according to the optimization evolution. Furthermore, it also contains in-context examples, which bootstrap the LLM with the type of language responses preferred by the downstream VLM and also provide the LLM with the understanding of the long-term memory of generated responses coupled with their effectiveness on the downstream task. Note that this prompt is used by our GLOV for CLIP. For the enhancing safety of VLMs, the objective is different, i.e., to reduce the ASR instead of increasing the accuracy.
@@ E.7 Safety Enhancement Prompts
4:  Generate $K$ prompts $P={List([}P_{1},P_{2},\dots,P_{K}{]})$.
@@
15:     NewPrompts $\leftarrow$ List([])